### v1

@@ Abstract
RGB-T tracking, a vital downstream task of object tracking, has made remarkable progress in recent years. Yet, it remains hindered by two major challenges: 1) the trade-off between performance and efficiency; 2) the scarcity of training data. To address the latter challenge, some recent methods employ prompts to fine-tune pre-trained RGB tracking models and leverage upstream knowledge in a parameter-efficient manner. However, these methods inadequately explore modality-independent patterns and disregard the dynamic reliability of different modalities in open scenarios. We propose M3PT, a novel RGB-T prompt tracking method that leverages middle fusion and multi-modal and multi-stage visual prompts to overcome these challenges. We pioneer the use of the middle fusion framework for RGB-T tracking, which achieves a balance between performance and efficiency. Furthermore, we incorporate the pre-trained RGB tracking model into the framework and utilize multiple flexible prompt strategies to adapt the pre-trained model to the comprehensive exploration of uni-modal patterns and the improved modeling of fusion-modal features, harnessing the potential of prompt learning in RGB-T tracking. Our method outperforms the state-of-the-art methods on four challenging benchmarks, while attaining 46.1 fps inference speed.
@@ 1 Introduction
Visual object tracking is a fundamental task in the computer vision community, which aims to continuously predict the state of a given target in every frame of a video sequence. It has a wide range of applications in fields such as autonomous driving [1], robot perception [2], and intelligent surveillance [3]. However, the tracking methods based on visible modality (also known as RGB tracking) are not robust in challenging scenarios such as extreme illumination and adverse weather, where the visible image quality is poor. To solve this problem, researchers have developed RGB-T tracking methods, which use the complementary information between thermal infrared and visible modalities to achieve all-weather and all-day tracking.
@@
RGB-T tracking data has a complex and rich cross-modal feature space, which is a correlated combination of the two modal feature spaces rather than a simple addition. Since the data of the two modalities are synchronized in time and space, they share multiple patterns such as object boundaries, partial fine-grained information, etc.; however, due to the different imaging bands, they also have many modality-specific and discriminative patterns. These patterns form the key information in the feature space, and exploring, integrating, and utilizing them is essential for robust RGB-T tracking. To achieve this goal, many excellent works [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15] have proposed various effective fusion-based multi-modal tracking frameworks. These frameworks can be classified into three categories based on the fusion location: image-level, feature-level, and decision-level fusion frameworks, as shown in Figure 1(a)-(c). The image-level fusion framework is simple and efficient, but it does not exploit the two modalities separately, leading to poor performance. On the other hand, the feature-level and decision-level fusion frameworks extract rich features from both modalities, achieving high performance, but they introduce redundant computations due to the dual-stream architecture, which reduces the inference efficiency. Therefore, a natural question arises: can we design a framework that optimizes both performance and efficiency?
@@
Figure 1: Comparison of different fusion tracking frameworks. (a)-(c) are the three mainstream frameworks: image-level, feature-level, and decision-level multi-modal fusion tracking frameworks, respectively. (d) is our proposed mid-term fusion tracking framework. Our framework, unlike the other three, splits the backbone into dual-stream and single-stream structures for uni-modal and fusion-modal feature representation, respectively, with the fusion module situated between them.
@@
Inspired by this idea, we develop a middle fusion framework for RGB-T tracking, as shown in Figure 1(d). We place the fusion module in the middle of the backbone, and split the backbone into two parts: the pre-fusion backbone extracts the information of the two modalities independently, and the post-fusion backbone enhances the fusion-modal features. By choosing an appropriate fusion location, this mixed-stream framework structure can achieve effective uni-modal modeling, as well as reduce the redundant feature extraction process, thus achieving the optimal balance between performance and efficiency. To the best of our knowledge, this is the first time that the middle fusion framework has been introduced into the RGB-T tracking task.
@@
Another major focus of current research is how to address the data scarcity problem. The high cost of data acquisition leads to a persistent lack of sufficient annotated training data for the RGB-T tracking task, which severely hinders the data-driven model from learning reliable task-specific knowledge. Since most RGB-T tracking methods are derived from upstream methods, some studies use pre-trained upstream models to initialize the parameters and facilitate faster and better convergence of the current model. However, these full fine-tuning methods require tuning a large number of parameters, which imposes extremely high demands on the computing and storage capabilities of the device, and may cause overfitting and forgetting of upstream knowledge.
@@
Considering these issues, some recent works [16, 17] introduce prompt learning into RGB-T tracking to better leverage the inheritance between upstream and downstream. With only tuning a small number of additional parameters, these methods can effectively reduce the gap between upstream and downstream tasks, and the training cost is much lower than the full fine-tuning methods. However, existing prompt-based methods have several key limitations: (i) Through direct or indirect image-level fusion, these methods hardly exploit any independent information within the modality, resulting in insufficient exploration and utilization of the discriminative patterns within the modality. (ii) They treat thermal infrared modality as an extra information only for generating prompts, and neglect the dynamic complementarity of the two modalities in open scenarios, thus failing to fully leverage the advantages of RGB-T tracking. (iii) They generate and inject prompts in a single manner.
@@
Motivated by the previous discussions, our goal is to employ adaptable visual prompts to fine-tune the pre-trained upstream model to identify and leverage the distinctive and supplementary information within multi-modal features, thus optimizing the performance of RGB-T prompt tracking. For this purpose, we introduce four innovative visual prompt strategies to transfer the upstream knowledge to the middle fusion tracking framework. Consequently, we devise an effective and efficient RGB-T prompt tracking method with middle fusion and multi-stage, multi-form prompts (M3PT). In the method, we divide the backbone of the foundation model into two parts, each for one of the two modeling stages of the middle fusion framework. To adapt the first part to uni-modal exploration, we design a Uni-modal Exploration Prompt Strategy, which uses the first part backbone and some lightweight prompters UEP to explore the modality-shared and modality-independent patterns layer by layer, and further generates reliable intra-modal and inter-modal prompts to guide the subsequent layers to model key features of the current modality and different modalities. Considering the dynamic complementarity of different modalities in open scenarios, we design a Middle Fusion Prompt Strategy to achieve adaptive selection and complementary fusion of the discriminative features of the two modalities, and use the fusion-modal features as natural prompts, which are fed into the second-stage backbone. To better adapt the backbone to the fusion-modal feature modeling, we design a Fusion-modal Enhancement Prompt Strategy, which guides the backbone to extract sufficient global and local fusion features, and thus obtain a richer fusion modality representation. Motivated by VPT, we design a Stage-aware and Modality-aware Prompt Strategy, which generates multiple learnable prompts to store the modality-fixed patterns and prepend them to the input of each stage. This strategy provides clear modality and stage indicators for the backbone. Notably, our method only contains 0.88M fine-tuning parameters. On four large-scale benchmarks, our method surpasses current state-of-the-art prompt-based methods by more than 2 percentage points on all metrics, and achieves competitive performance among existing methods. Moreover, our model has an inference speed of 46.1fps, exceeding most state-of-the-art methods. These results demonstrate the excellent balance of the proposed middle fusion framework in terms of performance and efficiency. Our work has the following main contributions:
@@
(1) We propose a middle fusion framework for RGB-T tracking task for the first time, which is flexible and achieves a good balance between performance and efficiency compared to previous RGB-T tracking frameworks.
@@
(2) Based on the proposed middle fusion framework, we develop a novel RGB-T prompt tracking method with multi-form and multi-stage visual prompts (M3PT), which fully unleashes the great potential of RGB-T prompt tracking through four flexible novel visual prompt strategies.
@@
(3) Extensive experiments on four challenging RGB-T tracking benchmarks validate the effectiveness, efficiency, and parameter-efficiency of our method.
@@ 2.1 RGB-T Tracking
Lu et al. [8] employ diverse adapters to capture the intra-modal and inter-modal relation of different modalities, and propose a new loss function that reduces the distribution divergence of multi-modal features across layers. Zhang et al. [9] develop a late fusion method that achieves robust multi-modal fusion through global and local weights, and uses Kalman filtering for motion estimation, to exploit the joint appearance and motion cues. Zhang et al. [10] extend the structure of siamese-like RGB trackers and introduce a weight generation model to enable the selection of unimodal discriminative features and the interaction of bimodal discriminative features. Tu et al. [7] propose a novel multi-modal multi-margin metric learning framework for RGBT tracking, which preserves the relations of multilevel hard samples during training. A major challenge for data-driven models in RGB-T tracking is the scarcity of labeled training data, which results from the high cost of data collection, alignment, and annotation. Compared to the abundant training data [32, 33, 34] for RGB tracking task, the training data [35, 15] for RGB-T tracking is severely inadequate for learning sufficient task-relevant knowledge.
@@
Meanwhile, some other works initialize the proposed RGB-T tracking models with RGB tracking models pre-trained on RGB tracking datasets, to guide the fast convergence of the models on limited training data. Zhang et al. [14] design a fusion module that can reduce the modality discrepancy and select the discriminative features, and initialize the RGB branch parameters with pre-trained DiMP parameters. Zhang et al. [15] integrate image-level, feature-level, and decision-level fusion into a unified framework, and use pre-trained DiMP model for parameter initialization. However, this approach requires a large number of parameters to be fine-tuned, which might cause the upstream knowledge to be forgotten during training, and also impose a heavy computational burden on the devices. To tackle this issue, some recent studies introduce the idea of prompt learning into RGB-T tracking, to achieve parameter-efficient fine-tuning and effective knowledge transfer from upstream tasks. We will introduce these methods in the next subsection.
@@ 2.2 Visual Prompt Learning
Prompt learning, as a parameter-efficient fine-tuning method, is originally proposed in the natural language processing field for transferring large language models to downstream tasks. Recently, this idea is introduced to the computer vision field [36, 37, 38, 39], using a small amount of image data from downstream tasks to generate effective visual prompts for transferring pre-trained models. Jia et al. [36] use a small number of learnable parameters as visual prompts, and prepend them to the input sequences of each layer. Bar et al. [37] propose to use the input-output pairs of downstream task examples and a new input image concatenated as visual prompts, to make the model learn to fill in the gaps, thus formulating this problem as an image inpainting problem. Considering that the downstream image datasets sometimes have large distribution differences, Huang et al. [38] cluster the image datasets into multiple subsets with the same internal distribution, configure unique prompts for each subset, and initialize them with a meta-prompt.
Yang et al. [16] take the weighted addition of two modal images as visual prompts and feed the prompts into the pre-trained RGB tracking model, which is the first time that prompt learning is introduced into the multi-modal tracking task. Zhu et al. [17] formulate the multi-modal tracking task as an RGB tracking task with an auxiliary modality input, and design a lightweight prompter to generate visual prompts from the auxiliary modality data. However, to bridge the gap between the upstream and downstream tasks, these methods overemphasize the role of the RGB modality, neglecting the dynamic dominance difference of different modalities in various scenarios, and thus fail to exploit modality-independent clues. In contrast, our method achieves effective exploration and adaptive fusion of different modality features by using more flexible and diverse prompt strategies, and therefore performs more robust when facing various challenging scenarios.
@@ 3 Preliminaries
As shown in Figure 2, we choose a one-stream one-stage RGB tracking model 2 as our foundation model. Built on ViT [40], the method is one-stream one-stage, which means that the feature extraction and template-search interaction are performed simultaneously by a single transformer backbone. The model consists of an embedding layer, a backbone composed of $L$ transformer encoder blocks stacked together, and a head network. The embedding layer maps the input template and search images into 16$\times$16 patches, reshapes them to $1D$ tokens and concatenates them as $E_{V}^{0}=[E_{V,Z}^{0};E_{V,X}^{0}]\in R^{(N_{Z}+N_{X})\times D}$, where $N_{Z}$ and $N_{X}$ respectively denote the number of template tokens and search tokens and $D$ denotes the dimension of the tokens. Then, the encoder blocks $\left\{Encoder^{l},l\in(1,L)\right\}$ perform self-attention mechanism [41] on the concatenated tokens $E_{V}^{0}$, thus achieving joint feature extraction and interaction between the template and the search:

### v2

@@ Abstract
RGB-T tracking, a vital downstream task of object tracking, has made remarkable progress in recent years. Yet, it remains hindered by two major challenges: 1) the trade-off between performance and efficiency; 2) the scarcity of training data. To address the latter challenge, some recent methods employ prompts to fine-tune pre-trained RGB tracking models and leverage upstream knowledge in a parameter-efficient manner. However, these methods inadequately explore modality-independent patterns and disregard the dynamic reliability of different modalities in open scenarios. We propose M3PT, a novel RGB-T prompt tracking method that leverages middle fusion and multi-modal and multi-stage visual prompts to overcome these challenges. We pioneer the use of the adjustable middle fusion meta-framework for RGB-T tracking, which could help the tracker balance the performance with efficiency, to meet various demands of application. Furthermore, based on the meta-framework, we utilize multiple flexible prompt strategies to adapt the pre-trained model to comprehensive exploration of uni-modal patterns and improved modeling of fusion-modal features in diverse modality-priority scenarios, harnessing the potential of prompt learning in RGB-T tracking. Evaluating on 6 existing challenging benchmarks, our method surpasses previous state-of-the-art prompt fine-tuning methods while maintaining great competitiveness against excellent full-parameter fine-tuning methods, with only 0.34M fine-tuned parameters.
@@ 1 Introduction
Visual object tracking is a fundamental task in the computer vision community, which aims to continuously predict the state of a given target in every frame of a video sequence. It has a wide range of applications in fields such as autonomous driving [1], robot perception [2], and intelligent surveillance [3]. However, the tracking methods based on visible modality (also known as RGB tracking) are not robust in challenging scenarios such as extreme illumination and adverse weather, where the visible image quality is poor. To solve this problem, researchers have developed RGB-T tracking methods [4, 5], which use the complementary information between thermal infrared and visible modalities to achieve all-weather and all-day tracking.
@@
RGB-T tracking data has a complex and rich cross-modal feature space, which is a correlated combination of the two modal feature spaces rather than a simple addition. Since the data of the two modalities are synchronized in time and space, they share multiple patterns such as object boundaries, partial fine-grained information, etc.; however, due to the different imaging bands, they also have many modality-specific and discriminative patterns. These patterns form the key information in the feature space, and exploring, integrating, and utilizing them is essential for robust RGB-T tracking. To achieve this goal, most excellent works [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23] are built on various effective fusion-based multi-modal tracking meta-frameworks. These meta-frameworks can be classified into three categories based on the fusion location: image-level, feature-level, and decision-level fusion frameworks, as shown in Figure 1(a)-(c). The image-level fusion framework is simple and efficient, but it does not exploit the two modalities separately, leading to poor performance. On the other hand, the feature-level and decision-level fusion frameworks extract rich features from both modalities, achieving high performance, but they introduce redundant computations due to the dual-stream architecture, which greatly reduces the inference efficiency. Therefore, a natural question arises: can we design a general RGB-T fusion tracking framework that enables trackers utilizing this framework to better balance efficiency with performance?
@@
Figure 1: Comparison of different fusion tracking frameworks. (a)-(c) are the three mainstream frameworks: image-level, feature-level, and decision-level multi-modal fusion tracking frameworks, respectively. (d) is our proposed middle fusion meta-framework. Our framework, unlike the other three, splits the backbone into dual-stream and single-stream structures for uni-modal and fusion-modal feature representation, respectively, with the fusion module situated between them.
@@
Inspired by this idea, we develop a general middle fusion meta-framework for RGB-T tracking, as shown in Figure 1(d). In this meta-framework, we place the fusion module in the middle of the backbone, and split the backbone into two parts: the pre-fusion backbone extracts the information of the two modalities independently, and the post-fusion backbone enhances the fusion-modal features. By choosing an appropriate fusion location and utilize effective fusion module, the trackers built on this mixed-stream meta-framework can achieve effective uni-modal modeling, as well as reduce the redundant feature extraction process, thus achieving a good balance between performance and efficiency. Also, by adjusting the fusion location, the efficiency and performance of the trackers could be changed flexibly for diverse demands. To the best of our knowledge, this is the first time that the middle fusion meta-framework has been introduced into the RGB-T tracking task.
@@
Another major focus of current research is how to address the data scarcity problem. The high cost of data acquisition leads to a persistent lack of sufficient annotated training data for the RGB-T tracking task, which severely hinders the data-driven model from learning reliable task-specific knowledge. Since most RGB-T tracking methods are derived from upstream methods, some studies [7, 17, 24, 23, 25, 22, 26] use pre-trained upstream models to initialize the parameters and facilitate faster and better convergence of the current model. However, these full fine-tuning methods require tuning a large number of parameters, which imposes extremely high demands on the computing and storage capabilities of the device, and may cause overfitting and forgetting of upstream knowledge.
@@
Considering these issues, some recent works [27, 28, 29, 30] introduce prompt learning into RGB-T tracking to better leverage the inheritance between upstream and downstream tasks. By tuning a minimal number of parameters, these methods effectively narrow the gap between upstream and downstream tasks, delivering competitive performance akin to most full fine-tuning approaches, while significantly reducing training costs. However, existing prompt-based methods have several key limitations: (i) Through direct or indirect image-level fusion, these methods hardly exploit the modality-independent information, resulting in insufficient exploration and utilization of the discriminative patterns within the modality. (ii) They treat thermal infrared modality as an extra information only for generating prompts and simply sum two modality features through constant global or element-wise weights, which neglect the dynamic modality-priority in intricate scenarios, thus failing to fully leverage the advantages of RGB-T tracking. (iii) They generate and inject prompts in a single manner like element-wise summation, which limits the great potential of prompt learning.
@@
Motivated by above discussions, we aim to employ adaptable visual prompts to fine-tune the pre-trained upstream model to identify and leverage the distinctive and supplementary information within multi-modal features, thus optimizing the performance of RGB-T prompt tracking. For this purpose, we introduce four innovative visual prompt strategies to transfer the upstream knowledge, based the middle fusion meta-framework. Consequently, we devise an novel and parameter-efficient RGB-T prompt tracking method with middle fusion and multi-stage, multi-form prompts (M3PT). In the method, we divide the backbone of the foundation model into two parts, each for one of the two modeling stages of the middle fusion framework. To adapt the first part to effective uni-modal and cross-modal modeling, we propose the Uni-modal and Inter-modal Exploration Prompt Strategy, which uses the first part backbone and two types of lightweight prompters UEP and IP to explore the modality-independent and cross-modal complementary patterns layer by layer, and further operate reliable intra-modal and inter-modal prompting to guide the subsequent layers to model key features of the current modality and different modalities. Considering the dynamic complementarity of different modalities in open scenarios, we design a Middle Fusion Prompt Strategy to achieve adaptive selection and complementary fusion of the discriminative features of the two modalities, and use the fusion-modal features as natural prompts, which are fed into the second-stage backbone. To better adapt the backbone to the fusion-modal feature modeling, we design a Fusion-modal Enhancement Prompt Strategy, which guides the backbone to extract sufficient global and local fusion features, and thus obtain a richer fusion modality representation. Motivated by VPT, we design a Stage-aware and Modality-aware Prompt Strategy, which generates multiple learnable prompts to store the modality-fixed patterns and prepend them to the input of each stage. This strategy provides clear modality and stage indicators for the backbone. Notably, our method only contains 0.34M fine-tuning parameters. On six challenging and large-scale benchmarks, our method surpasses all previous state-of-the-art prompt fine-tuning methods while maintaining great competitiveness against excellent full-parameter fine-tuning methods, with the least fine-tuned parameters. Our work has the following main contributions:
@@
(1) We propose a flexible and general middle fusion meta-framework for RGB-T tracking task for the first time. Building on this framework, the tracker could easily achieve a good balance between performance and efficiency, to meet diverse application demands.
@@
(2) Built on the proposed middle fusion meta-framework, we develop a novel RGB-T prompt tracking method with multi-form and multi-stage visual prompts (M3PT), which fully unleashes the great potential of RGB-T prompt tracking through four flexible novel visual prompt strategies.
@@
(3) Extensive experiments on six challenging RGB-T tracking benchmarks validate the effectiveness and parameter-efficiency of our method.
@@ 2.1 RGB-T Tracking
Lu et al. [10] employ diverse adapters to capture the intra-modal and inter-modal relation of different modalities, and propose a new loss function that reduces the distribution divergence of multi-modal features across layers. Zhang et al. [11] develop a late fusion method that achieves robust multi-modal fusion through global and local weights, and uses Kalman filtering for motion estimation, to exploit the joint appearance and motion cues. Zhang et al. [12] extend the structure of siamese-like RGB trackers and introduce a weight generation model to enable the selection of uni-modal discriminative features and the interaction of bimodal discriminative features. Tu et al. [9] propose a novel multi-modal multi-margin metric learning framework for RGB-T tracking, which preserves the relations of multilevel hard samples during training. A major challenge for data-driven models in RGB-T tracking is the scarcity of labeled training data, which results from the high cost of data collection, alignment, and annotation. Compared to the abundant training data [45, 46, 47] for RGB tracking task, the training data [17, 48] for RGB-T tracking is severely inadequate for learning sufficient task-relevant knowledge.
@@
Meanwhile, some recent works initialize the proposed RGB-T tracking models with RGB tracking models pre-trained on RGB tracking datasets. Zhang et al. [16] design a fusion module that can reduce the modality discrepancy and select the discriminative features, and initialize the RGB branch parameters with pre-trained DiMP parameters. Zhang et al. [17] integrate image-level, feature-level, and decision-level fusion into a unified framework, and use pre-trained DiMP model for parameter initialization. Based on pre-trained RGB siamese tracking parameters, Peng et al. [25] propose two subnetworks to fuse visible and infrared features and determine contributions of two modalities, respectively. Hui et al. [22] insert multiple template-bridged interaction modules into backbone to realize robust cross-modal interaction and utilize parameters from pre-trained transformer-based RGB tracking model for full fine-tuning. By employing full-parameter fine-tuning and a meticulously crafted cross-modal fusion module, these models achieve rapid convergence even with limited training data. Consequently, they exhibit significantly superior performance compared to other RGB-T trackers without full fine-tuning. However, the huge amount of tuned parameters also bring a heavy burden for training devices, in terms of computation and storage demands, while the potential forgetting of upstream knowledge should also be fixed attention. Considering above limitations of full fine-tuning methods, some contemporary researches integrate prompt learning into RGB-T tracking to facilitate streamlined parameter-efficient fine-tuning and the proficient transference of knowledge from upstream tasks. These methods will be introduced in the next subsection.
@@ 2.2 Visual Prompt Learning
Prompt learning, as a parameter-efficient transferring method, is originally proposed in the natural language processing field for transferring large language models to downstream tasks. Recently, this idea is introduced to the computer vision field [49, 50, 51, 52], using a small amount of image data from downstream tasks to generate effective visual prompts for transferring pre-trained models. Jia et al. [49] use a small number of learnable parameters as visual prompts, and prepend them to the input sequences of each layer. Bar et al. [50] propose to use the input-output pairs of downstream task examples and a new input image concatenated as visual prompts, to make the model learn to fill in the gaps, thus formulating this problem as an image inpainting problem. Considering that the downstream image datasets sometimes have large distribution differences, Huang et al. [51] cluster the image datasets into multiple subsets with the same internal distribution, configure unique prompts for each subset, and initialize them with a meta-prompt.
Yang et al. [27] take the weighted addition of two modal images as visual prompts and feed the prompts into the pre-trained RGB tracking model, which is the first time that prompt learning is introduced into the multi-modal tracking task. Zhu et al. [28] formulate the multi-modal tracking task as an RGB tracking task with an auxiliary modality input, and design a lightweight prompter to generate visual prompts from the auxiliary modality data. Hong et al. [30] further insert trainable adapters into foundation’s transformer layer for effective prompt fine-tuning. Wu et al. [29] combine three multimodal training sets and introduce lora techniques to learn common latent space of the additional modalities. However, to bridge the gap between the upstream and downstream tasks, these methods overemphasize the impact of the RGB modality, neglecting the dynamic modality-priority in various scenarios. Consequently, these methods fail to exploit modality-independent clues and struggle to perform robustly in intricate environments. Different with them, our prompt method effectively explores uni-modal information while adapting to scenarios through inter-modal interaction and fusion. This adaptability is achieved via flexible and diverse prompt strategies, enabling our method to excel when confronted with various challenging scenarios.
@@ 3 Preliminaries
As shown in Figure 2, we choose a one-stream one-stage RGB tracking model 2 as our foundation model. Built on ViT [53], the method is one-stream one-stage, which means that the feature extraction and template-search interaction are performed simultaneously by a single transformer backbone. The model consists of an embedding layer, a backbone composed of $L$ transformer encoder blocks stacked together, and a head network. The embedding layer maps the input template and search images into 16$\times$16 patches $E_{V,Z}^{0}\in R^{(N_{Z})\times D}$ and $E_{V,X}^{0}\in R^{(N_{Z})\times D}$, reshapes them to $1D$ tokens and concatenates them as $E_{V}^{0}=[E_{V,Z}^{0};E_{V,X}^{0}]\in R^{(N_{Z}+N_{X})\times D}$, where $N_{Z}$ and $N_{X}$ respectively denote the number of template tokens and search tokens and $D$ denotes the dimension of the tokens. Then, the encoder blocks $\left\{Encoder^{l},l\in(1,L)\right\}$ perform self-attention mechanism [54] on the concatenated tokens $E_{V}^{0}$, thus achieving joint feature extraction and interaction between the template and the search: